# Towards Imitation Learning in Real World Unstructured Social Mini-Games in Pedestrian Crowds

### v1

@@ Abstract
Imitation Learning (IL) strategies are used to generate policies for robot motion planning and navigation by learning from human trajectories. Recently, there has been a lot of excitement in applying IL in social interactions arising in urban environments such as university campuses, restaurants, grocery stores, and hospitals. However, obtaining numerous expert demonstrations in social settings might be expensive, risky, or even impossible. Current approaches therefore, focus only on simulated social interaction scenarios. This raises the question: How can a robot learn to imitate an expert demonstrator from real world multi-agent social interaction scenarios? It remains unknown which, if any, IL methods perform well and what assumptions they require.
We benchmark representative IL methods in real world social interaction scenarios on a motion planning task, using a novel pedestrian intersection dataset collected at the University of Texas at Austin campus. Our evaluation reveals two key findings: first, learning multi-agent cost functions is required for learning the diverse behavior modes of agents in tightly coupled interactions and second, conditioning the training of IL methods on partial state information or providing global information in simulation can improve imitation learning, especially in real world social interaction scenarios.
@@ I Introduction
Robot navigation in densely populated human environments has garnered significant attention. Of particular interest are tightly coupled social interactions between pedestrians, scooters, and vehicles in geometrically constrained scenarios such as intersections, hallways, narrow gaps, etc. These scenarios, which we refer to as Social Mini-Games or SMGs [1], occur both in outdoor settings [2, 3, 4, 5] as well as indoor spaces like restaurants, grocery stores, hospitals, and university campuses [6, 7, 8, 9, 10, 11]. Despite some success, deploying robots in SMGs in a socially compliant manner, such as yielding to a pedestrian, is challenging as social compliance requires predicting and reacting to humans’ intents as well as projecting their own [12, 13]. Since humans are most adept at navigating SMGs in a socially compliant way, imitation learning (IL) has been instrumental in social robot navigation [6] where robot navigation policies are learned by mimicking human trajectories. However, acquiring large volumes of expert demonstration data in urban settings is not only resource-intensive but can also be challenging due to safety and privacy concerns. This constraint has limited the state-of-the-art approaches to work primarily in simulated environments leading us to ask: Can imitation learning be used to learn robot navigation policies for real world SMGs?
Figure 1: Visualizing the imitation learning problem for motion planning at the Speedway intersection on the University of Texas at Austin campus. We refer to such interactions as social mini-games.
This research sets out to bridge this knowledge gap by studying the strengths, limitations, and the various assumptions made by various IL methods in real world SMGs. Using a novel dataset collected on the University of Texas at Austin campus, we benchmark and evaluate the comparative performance of these methods in pedestrian-rich contexts. Through this study, we aim to guide the selection of appropriate IL techniques for motion planning in large pedestrian crowds, ensuring that robots can navigate shared spaces safely and with social awareness.
The methods benchmarked in this study follow the taxonomy laid out in seminal surveys on imitation learning [14, 15] that split IL methods broadly into behavior cloning (BC), inverse reinforcement learning (IRL), and generative methods. We pick, and compare, representative methods from each class. For instance, the representative method for IRL is the original Maximum Entropy IRL (MaxEnt IRL) algorithm proposed by Ziebart et al. [16].
Main contributions: In this work,
  1. 1.
    We present the first study of IL methods, such as behavior cloning, generative models, and inverse reinforcement learning, applied to motion planning in real pedestrian crowds. Our evaluation reveals two key findings: $(i)$, learning multi-agent cost functions is required for learning the diverse behavior modes of agents in tightly coupled interactions and $(ii)$ conditioning the training of behavior cloning on partial state information or providing global information in simulation can improve imitation learning, especially in real world social interaction scenarios.
  2. 2.
    We release a new dataset [^1] https://amrl.cs.utexas.edu/imitation-learning-benchmark.html collected on the campus of the University of Texas at Austin. It consists of $3$ hours of trajectory data collected at a busy intersection consisting of positions, velocity, heading angle, and class ids of pedestrians. Our dataset primarily consists of tightly coupled agent interactions that are rarely found in existing autonomous driving or urban pedestrian datasets.
II Imitation Learning: Background
IL represents a machine learning paradigm where an autonomous agent strives to learn a behavior by emulating an expert’s demonstrations. Typically, the demonstrations comprise state-action pairs, generated by an expert policy during real-world execution. Addressing the imitation learning problem has led to three major approaches, namely, Behavior Cloning (BC) [17, 18, 19], which involves direct learning of an imitative policy through supervised learning, Inverse Reinforcement Learning (IRL) [20, 21, 22], wherein a reward function is first inferred from the demonstrations, subsequently used to guide policy learning via Reinforcement Learning (RL) [23], and generative models [24, 25, 26]. We refer the reader to Zhang et al. [15] for a more comprehensive background on IL.
II-A Behavior Cloning
Given a sequence of tuples containing state-action pairs in the expert demonstration, Behavior Cloning (bc) employs a supervised learning approach to learn the imitator’s policy mapping from states to actions [27, 28]. Despite its relative ease of implementation and ability to scale and perform effectively in a large-scale data regime, bc is known to suffer from a compounding error issue [29]. This problem typically arises due to a slight deviation from the expert trajectory leading to drastic errors over time, a characteristic feature of bc. Regardless, the primary focus of this work is on offline imitation learning techniques, which are uniquely designed to operate with a static, comprehensive dataset of expert demonstrations. These methods hold considerable promise for enhancing efficiency and reproducibility in complex tasks, particularly in multi-agent human motion prediction.
II-B Inverse Reinforcement Learning
While traditional IRL methods achieved success, they were primarily designed for discrete state and action spaces. In contrast, Levine et al. investigated maximum entropy cost inference for systems with continuous state and action spaces [30]. A common assumption in these studies involved parameterizing the agent’s cost function as a weighted sum of known features. This assumption was relaxed in Finn et al.’s work [31], where the underlying cost function was learned using a neural network within the maximum-entropy framework. In the context of bilevel optimization, works like [32, 33] focused on inferring an agent’s cost function by optimizing the estimation error of system trajectories or the likelihood of demonstrations.
@@ II-B Inverse Reinforcement Learning
While previous approaches focused on learning in the single-agent regime, multi-Agent IRL has been extensively studied in various settings. Adversarial machine learning techniques were employed for high-dimensional state and action spaces [34], and generative adversarial imitation learning was extended to the multi-agent setting [35]. Additionally, the problem has been approached as an estimation task [36, 37], and linear quadratic games were considered with known equilibrium strategies [38, 39]. The maximum entropy IRL [40] framework was introduced to accommodate bounded rationality and noisy human demonstrations. These studies highlight the diverse and evolving landscape of multi-agent IRL.
II-C Generative Models
In a recent development, a novel category of imitation learning algorithms known as Adversarial Imitation Learning (AIL) [24, 25, 26] has emerged. AIL introduces an intriguing adversarial setup akin to the concept of Generative Adversarial Networks (GANs). Here, the agent employs two neural networks: a discriminator, responsible for distinguishing expert experiences from the imitator’s, and a generator network that models the imitator’s policy, mimicking the expert demonstrations. Among the pioneering AIL algorithms stands out Generative Adversarial Imitation Learning (GAIL) [24], showcasing remarkable ability in replicating expert behavior. GAIL directly extracts a policy from an agent’s demonstrations, similar to reinforcement learning following maximum entropy IRL.
However, a shared drawback among these algorithms is their underlying assumption of having access to the expert’s action data during the demonstration phase, which may not be readily available in all practical scenarios [41]. Additionally, these methods learn from a single agent in isolation.
III Problem Formulation
@@
We define a game, $G\coloneqq\left\langle k,\mathcal{X},\mathcal{T},\{\mathcal{U}^{i}\},\{\mathcal{J}^{i}\}\right\rangle$, where $k$ denotes the number of agents. Hereafter, $i$ will refer to the index of an agent and appear as a superscript whereas $t$ will refer to the current time-step and appear as a subscript. The general state space $\mathcal{X}$ (e.g. SE(2), SE(3), etc.) is continuous; the $i^{\textrm{th}}$ agent at time $t$ has a state $x^{i}_{t}\in\mathcal{X}^{i}$. Over a finite horizon $T$, each agent starts from $x^{i}_{0}\in\mathcal{X}^{i}$ and reaches a goal state $x^{i}_{T}\in\mathcal{X}^{i}$. An agent’s transition function is a mapping, $\mathcal{T}^{i}:\mathcal{X}\times\mathcal{U}^{i}\longrightarrow\mathcal{X}$, where $u^{i}_{t}\in\mathcal{U}^{i}$ is the continuous control space for agent $i$. A discrete trajectory is specified by the sequence $\Gamma^{i}=\left(x^{i}_{0},u^{i}_{0},x^{i}_{1},u^{i}_{1},\ldots,x^{i}_{T-1},u^{i}_{T-1},x^{i}_{T}\right)$. Each agent has a parameterized running cost $\mathcal{J}^{i}:\mathcal{X}\times\mathcal{\mathbf{U}}\longrightarrow\mathbb{R}$ where $\mathcal{\mathbf{U}}=\mathcal{U}^{1}\times\mathcal{U}^{2}\times,\ldots,\times\mathcal{U}^{k}$. The cost $\mathcal{J}^{i}$ acts on joint state $x_{t}$ and control $\mathbf{u}_{t}\in\mathbf{U}$ at each time step measuring the distances between agents, deviation from the reference path, and the change in the control input. Each agent $i$ follows a stochastic decentralized policy $\mathcal{F}^{i}_{\theta^{i}}:\mathcal{X}\longrightarrow\mathcal{U}^{i}$, which captures the probability of agent $i$ selecting action $u^{i}_{t}$ at time $t$, given that the system is in state $x_{t}$. The parameters for the policy $\mathcal{F}^{i}$ are represented by $\theta^{i}$, which may correspond to weights of a neural network or coefficients in a weighted linear function. We assume that we are provided with a dataset $\mathcal{\widehat{D}}=\{\widehat{\Gamma}^{1:k}\}_{N}$ that consists of $N$ expert demonstration trajectories, $\widehat{\Gamma}^{i}$ for each agent $i$. Further, we let $\mathcal{D}_{\theta^{i}}=\{\Gamma^{1:k}_{\theta^{i}}\}_{M}$ denote the dataset of $M$ trajectories for each agent generated by the policy $\mathcal{F}^{i}_{\theta^{i}}$ parameterized by $\theta^{i}$ (in the case of IRL, $\mathcal{F}^{i}_{\theta^{i}}$ implies the reward function is parameterized by $\theta^{i}$). Our problem statement is then,
@@
where $\Psi^{i}\left(\Gamma^{i}_{\theta^{i}},\widehat{\Gamma}^{i}\right)$ is a generalized discrepancy function that takes on different forms depending on the problem formulation (IRL versus BC, etc.). For instance,
$\Psi^{i}\left(\Gamma^{i}_{\theta^{i}},\widehat{\Gamma}^{i}\right)=\begin{cases}\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)-\Phi^{i}\left(\widehat{\Gamma}^{i}\right)&\text{for IRL}\\ \frac{1}{T}\sum_{t=0}^{T-1}\mathcal{F}^{i}_{\theta^{i}}\left(x^{i}_{t}\right)-u^{i}_{t}&\text{for BC}\end{cases}$
where $\Phi(\cdot)$ is a feature function that, given a trajectory, returns a vector of features $\left[\phi_{1},\phi_{2},\ldots,\phi_{m}\right]^{\top}$, averaged across the length of the trajectory, where $m$ refers to the number of distinct features. A simple example of a feature function is the weighted linear sum $\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)=\sum_{t=0}^{T-1}{\theta^{i}}^{\top}\mathcal{J}^{i}\left(x_{t},\mathbf{u}_{t}\right)$. In IRL, we want to find parameters $\theta^{i}$ for each agent $i$ such that the feature expectation of the generated trajectories $\widehat{\Gamma}^{i}$ under the learned cost weights is as close to the empirical feature expectation of the expert demonstrations. Note that we slightly abuse notation and ignore terminal cost at $t=T$.
IV IL Approaches for Pedestrian Crowds
Representative BaselinesIRLBehavior CloningGenerative ModelsHybrid IRL-BCMulti-agent MaxEnt IRLMaxEnt IRLMLP-basedLSTM-basedTransformer-basedMLP-basedLSTM-basedTransformer-basedCVAE-basedImplicit behavior cloning
Figure 2: Representative baselines used for evaluation and benchmarking. The boxed text represents the different algorithms and the underlined text represents the different architectures. Methods listed under the algorithms and architectures indicate usage of the corresponding algorithm or architecture, or both.
In this section, we discuss the technical approaches underpinning the methods included in our evaluation. Our selection criteria emphasizes both breadth, in terms of different algorithms used and depth, in terms of representing a large set of contemporary approaches. We visually summarize the selection of baselines in Figure 2.
Following the classification of imitation learning in the literature, we discuss behavior cloning (BC), inverse reinforcement learning (IRL), and generative models. Any hybrid models are represented as a separate category. We define a representative algorithm as the best performing method in its category as reported in the literature. For instance, PECNet [42] is the top performing generative method using a conditional variational autoencoder on the ETH/UCY trajectory forecasting benchmark, therefore, it is treated as a representative method for that category.
IV-A Inverse Reinforcement Learning [43, 40]
Maximum Entropy IRL [43]. In this section, we describe the basic Maximum Entropy Inverse Reinforcement Learning (MaxEnt IRL) approach tailored for deriving reward functions from observed expert trajectories [43, 16], followed by highlighting its benefits and the main reasons for its inclusion as a representative baseline.
Traditional Inverse Reinforcement Learning (IRL) seeks to infer a cost function from expert demonstrations [27]. However, multiple cost functions can rationalize the expert’s behavior, leading to inherent ambiguities. MaxEnt IRL addresses this by incorporating the principle of maximum entropy, ensuring that the derived cost function is both unique and maximally uncertain. Given expert trajectories, the probability of a trajectory $\Gamma^{i}$ under the MaxEnt framework for agent $i$ is defined as:

### v2

# Multi-Agent Inverse Reinforcement Learning in Real World Unstructured Pedestrian Crowds

@@ Abstract
Social robot navigation in crowded public spaces such as university campuses, restaurants, grocery stores, and hospitals, is an increasingly important area of research. One of the core strategies for achieving this goal is to understand humans’ intent–underlying psychological factors that govern their motion–by learning their reward functions, typically via inverse reinforcement learning (IRL). Despite significant progress in IRL, learning reward functions of multiple agents simultaneously in dense unstructured pedestrian crowds has remained intractable due to the nature of the tightly coupled social interactions that occur in these scenarios e.g. passing, intersections, swerving, weaving, etc. In this paper, we present a new multi-agent maximum entropy inverse reinforcement learning algorithm for real world unstructured pedestrian crowds. Key to our approach is a simple, but effective, mathematical trick which we name the so-called “tractability-rationality trade-off” trick that achieves tractability at the cost of a slight reduction in accuracy. We compare our approach to the classical single-agent MaxEnt IRL as well as state-of-the-art trajectory prediction methods on several datasets including the ETH, UCY, SCAND, JRDB, and a new dataset, called Speedway, collected at a busy intersection on a University campus focusing on dense, complex agent interactions. Our key findings show that, on the dense Speedway dataset, our approach ranks $1^{\textrm{st}}$ among top $7$ baselines with >2×>2$\times$> 2 × improvement over single-agent IRL, and is competitive with state-of-the-art large transformer-based encoder-decoder models on sparser datasets such as ETH/UCY (ranks $3^{\textrm{rd}}$ among top $7$ baselines).
@@ I Introduction
Robot navigation in densely populated human environments has garnered significant attention, especially in scenarios involving tightly coupled social interactions among pedestrians, scooters, and vehicles. These scenarios [1] occur in both outdoor settings [2, 3, 4, 5] and indoor spaces such as restaurants, grocery stores, hospitals, and university campuses [6, 7, 8, 9, 10, 11]. Effective navigation in such dense and cluttered human environments is challenging requires robots to be socially compliant which involves predicting and responding to human intentions [12, 13]. Many approaches have approached intent prediction implicitly via trajectory forecasting [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. While trajectory forecasting has yielded impressive results on sparse crowd datasets such as ETH, UCY, and JRDB, and structured traffic datasets such as the WAYMO Motion Forecasting Dataset [25], it is unclear whether they perform similarly well in more complex settings that include passing, weaving, yielding to others, etc.
Alternatively, some approaches have modeled intent explicitly by computing humans’ reward functions via inverse reinforcement learning. Recently, Gonon and Billard [26] used maximum entropy IRL [27] for achieving socially compliant navigation in pedestrian crowds. However, inherently a single-agent approach, maximum entropy IRL (MaxEnt IRL) assumes a single objective function for all the agents, which works reasonably well in sparse crowds that lack many tightly coupled interactions, as shown in [26], but fails to accommodate humans’ objectives in denser and more unstructured pedestrian crowds, as we show in our experiments in Section V. This limitation leads us to question: Can MaxEnt IRL be adapted to learn effective robot navigation policies for real-world multi-agent unstructured pedestrian crowds?
Figure 1: Quantitatively benchmarking representative IL approaches on the Speedway dataset: behavior cloning (BC [28]), inverse reinforcement learning (IRL [27] and MAIRL [29]), and generative methods (PECNet [30]). (left) Root mean square error per timestep; larger slope indicates poor long-term planning. (right) Cumulative RMSE distribution demonstrates the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective learner.
The problem might suggest a multi-agent IRL (Inverse Reinforcement Learning) approach, for which numerous solutions already exist in the dynamic game theory and reinforcement learning literature [31, 32, 33, 34, 35, 36, 37, 38, 29]. The primary limitation to multi-agent IRL is that it assumes known dynamics/system models of all the agents in the environment and the availability of large amounts of training data.
Main contribution: In this work, we propose a novel multi-agent MaxEnt IRL algorithm for learning reward functions of humans in real world unstructured pedestrian crowds. The key insight in our algorithm is a simple, but effective, mathematical trick which we name the so-called “tractability-rationality trade-off” trick that essentially relaxes the assumption of known system dynamics at the cost of a slight reduction in accuracy. Notably, this trade-off not only makes multi-agent IRL tractable in unstructured settings but also paves the way for new research directions—such as computing exact bounds for this trade-off.
We use off-the-shelf dynamic game theory solvers to roll out trajectories that maximize the learned reward functions. We evaluate this new approach using a dataset collected at a busy intersection on a University campus comprising of pedestrians, scooters, bicycles, particularly focusing on tightly coupled interactions such as when agents walk towards each other, walk past each other, squeeze past each other, etc.We compare our approach to the classical single-agent MaxEnt IRL as well as state-of-the-art trajectory prediction methods on several datasets including the ETH, UCY, JRDB, and a new dataset collected at a busy intersection on a University campus focusing on complex agent interaction. Our findings highlight two key insights: $(i)$ our approach yields the lowest error between the policy-generated trajectories and the ground truth trajectories, $(ii)$ our approach has the highest percentage of low-RMSE trajectories.
II Related Work
II-A Inverse Reinforcement Learning
Kalman initially explored the inference of an agent’s cost function within the context of inverse optimal control for linear quadratic systems with a linear control law [39]. Subsequent works [40, 41] considered learning the agent’s cost, assuming demonstrations adhered to optimality conditions. However, multiple cost functions can rationalize the expert’s behavior, leading to inherent ambiguities. MaxEnt IRL [27] addresses this by incorporating the principle of maximum entropy, ensuring that the derived cost function is both unique and maximally uncertain.
@@ II-A Inverse Reinforcement Learning
Multi-agent IRL: While previous approaches focused on learning in the single-agent regime, multi-agent IRL has been extensively studied in various settings. Initial efforts focused on discrete state and action spaces [31], while later research extended to general-sum games with discrete spaces [32]. Adversarial machine learning techniques were employed for high-dimensional state and action spaces [33], and generative adversarial imitation learning was extended to the multi-agent setting [34]. Additionally, the problem has been approached as an estimation task [35, 36], and linear quadratic games were considered with known equilibrium strategies [37, 38]. The maximum entropy IRL [29] framework was introduced to accommodate bounded rationality and noisy human demonstrations.
II-B Trajectory Prediction
We focus on deep learning-based trajectory and intent prediction which dates back ten years to the seminal SocialLSTM work [42]. TraPHic and RobustTP [43, 44] use an LSTM-CNN framework to predict trajectories in dense and complex environments. TNT [45] uses target prediction, motion estimation, and ranking-based trajectory selection to predict future trajectories. DESIRE [46] uses sample generation and trajectory ranking for trajectory prediction. PRECOG [47] combines conditioned trajectory forecasting with planning objectives for agents. [48] uses a Seq2Seq framework to encode agents’ observations over neighboring agents as their social context for trajectory forecasting and decision-making. [49] uses temporal smoothness in attention modeling for interactions and a sequential model for trajectory prediction. However, a major limitation of this class of methods is that they employ behavior cloning to try and mimic the data, without inherently capturing humans’ internal reward functions.
II-C Intent-aware Multi-agent Reinforcement Learning
Related to our task, intent-aware multi-agent reinforcement learning [50] estimates an intrinsic value that represents opponents’ intentions for communication [51] or decision-making. Many intent inference modules are based on Theory of Mind (ToM) [52] reasoning or social attention-related mechanisms [53, 54]. [55] uses ToM reasoning over opponents’ reward functions from their historical behaviors in performing multi-agent inverse reinforcement learning (MAIRL). [9] uses game theory ideas to reason about other agents’ incentives and help decentralized planning among strategic agents. However, many prior works oversimplify the intent inference and make some prior assumptions about the content of intent.
II-D Opponent Modeling
Opponent modeling [56] has been studied in multi-agent reinforcement learning and usually deploys various inference mechanisms to understand and predict other agents’ policies. Opponent modeling could be done by either estimating others’ actions and safety via Gaussian Process [57] or by generating embeddings representing opponents’ observations and actions [58]. Inferring opponents’ policies helps to interpret peer agents’ actions [59] and makes agents more adaptive when encountering new partners [60]. Notably, many works [61, 62] reveal the phenomenon whereby ego agents’ policies also influence opponents’ policies. To track the dynamic variation of opponents’ strategies made by an ego agent’s influence, [63, 64] propose the latent representation to model opponents’ strategies and influence based on their findings on the underlying structure in agents’ strategy space. [65] provides a causal influence mechanism over opponents’ actions and defines an influential reward over actions with high influence over others’ policies. [66] proposes an optimization objective that accounts for the long-term impact of ego agents’ behavior in opponent modeling. A considerable limitation of many current me thodologies is the underlying assumption that agents continually interact with a consistent set of opponents across episodes.
Placing our work in the literature: Our work directly builds on top of the multi-agent IRL literature. It is complementary to the trajectory prediction, multi-agent reinforcement learning, and opponent modeling literature, which implicitly model other agent’s objectives, whereas our approach explicitly models humans’ intent via reward functions.
III Problem Formulation and Background
[…]
If $\left(H^{i}_{t}\right)^{-1}$ is non-PSD, perform the Tractability-Rationality Trade-off trick (18)
Until convergence $\theta^{i}_{\textrm{new}}$ $\Gamma^{i}_{\theta^{i}_{\textrm{new}}}$
Figure 2: Flow diagram of our multi-agent inverse reinforcement learning algorithm. The algorithm begins by taking as input the system dynamics, initial state, dataset, feature function, and initial parameters. At each time step, it computes the expected features based on the current policy. These features are used to update the weights by minimizing the difference between the expected and ground truth features. The updated weights are then used to update the game-theoretic policy and roll out a new set of trajectories. If the Hessian matrix is non-positive definite, the ”tractability-rationality trade-off” trick is applied to ensure tractability in unstructured environments.
@@ III Problem Formulation and Background
We define a game, $G\coloneqq\left\langle k,\mathcal{X},\mathcal{T},\{\mathcal{U}^{i}\},\{\mathcal{J}^{i}\}\right\rangle$, where $k$ denotes the number of agents. Hereafter, $i$ will refer to the index of an agent and appear as a superscript whereas $t$ will refer to the current time-step and appear as a subscript. The general state space $\mathcal{X}$ (e.g. SE(2), SE(3), etc.) is continuous; the $i^{\textrm{th}}$ agent at time $t$ has a state $x^{i}_{t}\in\mathcal{X}^{i}$. Over a finite horizon $T$, each agent starts from $x^{i}_{0}\in\mathcal{X}^{i}$ and reaches a goal state $x^{i}_{T}\in\mathcal{X}^{i}$. An agent’s transition function is a mapping, $\mathcal{T}^{i}:\mathcal{X}\times\mathcal{U}^{i}\longrightarrow\mathcal{X}$, where $u^{i}_{t}\in\mathcal{U}^{i}$ is the continuous control space for agent $i$. A discrete trajectory is specified by the sequence $\Gamma^{i}=\left(x^{i}_{0},u^{i}_{0},x^{i}_{1},u^{i}_{1},\ldots,x^{i}_{T-1},u^{i}_{T-1},x^{i}_{T}\right)$.
Each agent has a parameterized running cost [^1] Note that broadly speaking, reward maximization is equivalent to cost minimization, which is what we model in this work. $\mathcal{J}^{i}:\mathcal{X}\times\mathcal{\mathbf{U}}\longrightarrow\mathbb{R}$ where $\mathcal{\mathbf{U}}=\mathcal{U}^{1}\times\mathcal{U}^{2}\times,\ldots,\times\mathcal{U}^{k}$. The cost $\mathcal{J}^{i}$ acts on joint state $x_{t}$ and control $\mathbf{u}_{t}\in\mathbf{U}$ at each time step measuring the distances between agents, deviation from the reference path, and the change in the control input. Each agent $i$ follows a stochastic decentralized policy $\mathcal{F}^{i}_{\theta^{i}}:\mathcal{X}\longrightarrow\mathcal{U}^{i}$, which captures the probability of agent $i$ selecting action $u^{i}_{t}$ at time $t$, given that the system is in state $x_{t}$. The parameters for the policy $\mathcal{F}^{i}$ are represented by $\theta^{i}$, which may correspond to weights of a neural network or coefficients in a weighted linear function. We assume that we are provided with a dataset $\mathcal{\widehat{D}}=\{\widehat{\Gamma}^{1:k}\}_{N}$ that consists of $N$ expert demonstration trajectories, $\widehat{\Gamma}^{i}$ for each agent $i$. Further, we let $\mathcal{D}_{\theta^{i}}=\{\Gamma^{1:k}_{\theta^{i}}\}_{M}$ denote the dataset of $M$ trajectories for each agent generated by the policy $\mathcal{F}^{i}_{\theta^{i}}$ parameterized by $\theta^{i}$ (in the case of IRL, $\mathcal{F}^{i}_{\theta^{i}}$ implies the reward function is parameterized by $\theta^{i}$). Our problem statement is then,
@@
where $\Psi^{i}\left(\Gamma^{i}_{\theta^{i}},\widehat{\Gamma}^{i}\right)$ is a generalized discrepancy function,
$\Psi^{i}\left(\Gamma^{i}_{\theta^{i}},\widehat{\Gamma}^{i}\right)=\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)-\Phi^{i}\left(\widehat{\Gamma}^{i}\right)$ (2)
and $\Phi(\cdot)$ is a feature function that, given a trajectory, returns a vector of features $\left[\phi_{1},\phi_{2},\ldots,\phi_{m}\right]^{\top}$, averaged across the length of the trajectory, where $m$ refers to the number of distinct features. A simple example of a feature function is the weighted linear sum $\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)=\sum_{t=0}^{T-1}{\theta^{i}}^{\top}\mathcal{J}^{i}\left(x_{t},\mathbf{u}_{t}\right)$. In IRL, we want to find parameters $\theta^{i}$ for each agent $i$ such that the feature expectation of the generated trajectories $\widehat{\Gamma}^{i}$ under the learned cost weights is as close to the empirical feature expectation of the expert demonstrations. Note that we slightly abuse notation and ignore terminal cost at $t=T$.
III-A Background: MaxEnt IRL
As background, we first briefly review the key details of MaxEnt IRL [26, 27]. Given expert trajectories, the probability of a trajectory $\Gamma^{i}$ under the MaxEnt framework for agent $i$ is defined as: